3/8 = 15/40 3/8 = 15/40 3/8 = 15/40 3/8 = 15/40 3/8 = 15/40

 3/8 = 15/40
 3/8 = 15/40
 3/8 = 15/40
 3/8 = 15/40
 3/8 = 15/40

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is True.

Consider the two fractions, 3/8 = 15/40
15/40 is further simplified by dividing both numerator and denominator by 5
we get, = 3/8
Therefore, 3/8 = 3/8

Thanks